# Context Representation via Action-Free Transformer encoder-decoder for Meta Reinforcement Learning

### v1

@@ Abstract
Reinforcement learning (RL) enables robots to operate in uncertain environments, but standard approaches often struggle with poor generalization to unseen tasks. Context-adaptive meta reinforcement learning addresses these limitations by conditioning on task representation, yet most methods rely on full trajectories, including actions, making task inference tightly coupled to a specific policy. This paper introduces Context Representation via an Action-Free Transformer encoder–decoder (CRAFT), a belief model that infers task representations solely from sequences of states and rewards. By removing the dependence on actions, CRAFT decouples task inference from policy optimization, supports modular training, and leverages amortized variational inference for scalable belief updates. Built on a transformer encoder–decoder with rotary positional embeddings, the model captures long-range temporal dependencies and robustly encodes both parametric and non-parametric task variations. Experiments on the MetaWorld ML-10 robotic manipulation benchmark show that CRAFT achieves faster adaptation, improved generalization, and more effective exploration compared to context-adaptive meta-RL baselines. These findings highlight the potential of action-free inference as a foundation for scalable RL in robotic control. [^1] This work has been published in part in the doctoral dissertation of Amir M. Soufi Enayati [enayati2025practical].
@@ 1 Introduction
Figure 1: An example of parametric and non-parametric variations in meta-RL
Although Bayes-optimal agents offer ideal exploration-exploitation trade-offs, they are intractable in complex domains. Posterior sampling and latent variable meta-learning attempt to approximate this behavior, encoding task uncertainty into structured representations to reduce reliance on extensive interactions. Yet, most meta-RL approaches struggle with significant task variability, often assuming smooth parametric changes between tasks, which limits their adaptability in real-world robotic scenarios involving non-parametric variations [zintgraf2019varibad]. A schematic of different approaches in context accumulation for exploration in meta-RL is depicted in Figure˜2.
Figure 2: Exploration approaches to meta-RL: posterior sampling (left), Bayes-optimal (middle), Bayes-adaptive (right)
@@
To address current limitations, a framework is proposed in this paper for extracting long-term interepisodic features for Bayes-adaptive exploration. Also, CRAFT isolates the agent’s performance from task history by removing the condition of the belief model on the actions. This abstraction will allow learned behaviors to be reused and recombined in every stage of the training. Leveraging structured latent spaces and variational inference, the method facilitates strategic exploration and robust generalization, evaluated in a robotic simulation benchmark. This integration of probabilistic task embeddings and meta-learning techniques aims to ameliorate performance limitations and advance autonomy in RL-driven robot manipulation.
@@ 2.1 Meta-RL General Outline
MAML is a gradient-based optimization problem to formulate proximity after a few adaptation steps. With the loss function for each task $\mathcal{T}_{\mathrm{i}}$ defined as
$\mathcal{L}_{\mathcal{T}_{\mathrm{i}}}(\pi_{\theta})=-\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta}}\left[\sum_{t=0}^{N_{\tau}}\gamma^{t}r_{t+1})\right],$ (1)
@@
Another challenge in meta-RL is the problem of generalization across tasks and domains. Meta-RL algorithms aim to learn meta-policies that can adapt to new tasks or environments with minimal additional training. However, achieving robust generalization remains a significant issue. Meta-learned policies often struggle to transfer knowledge to tasks or domains that differ significantly from the training distribution, resulting in poor performance and limited applicability. yu2020meta has shown that state-of-the-art meta-RL algorithms usually struggle in adaptation to an unseen robotic manipulation task after being exposed to ten or even 45 tasks in meta-training. This indicates that knowledge sharing among the training tasks is not easily generalizable to the test domain.
@@ 2.2 Bayesian Approach To Context-Adaptive Meta-RL
$=\mathbb{E}_{b_{t}}\left[\mathcal{P}_{S}(s_{t+1}|s_{t},a_{t+1})\right]~\delta(b_{t+1}=p(\mathcal{P}_{R},\mathcal{P}_{S}|\tau_{0:t+1})).$ (6)
@@ 3.2 The Action-Free Belief Model Design
$\mathcal{P}_{R}(r|s,a,s^{\prime})\approx\tilde{\mathcal{P}}_{R}(r|s_{0},\dots,s_{-1},s,s^{\prime})\quad\text{and}\quad\mathcal{P}_{S}(s^{\prime}|s,a)\approx\tilde{\mathcal{P}}_{S}(s^{\prime}|s_{0},\dots,s_{-1},s).$
@@ 3.3.1 Rotary Positional Embedding
where $i$ is the token position in the context and $t\in{0,\dots,\frac{d}{2}-1}$ with $d$ representing the length of the token.
@@ 3.3.2 Implementation in CRAFT
The architecture shown in Figure˜3(a) uses a transformer encoder-decoder to extract compact and informative representations from action-free trajectories, aligning with the structure of belief updates in Bayes-Adaptive Markov Decision Processes (BAMDP). Borrowing from Equation˜12, the posterior belief over task identity is defined as
@@
$[\mathbf{u}^{s}_{i}]_{i=0:t+1}$ $=\text{Causal SelfAttn}(\mathbf{q},\mathbf{k},\mathbf{v})$
$=\text{Causal SelfAttn}([f_{q}(\mathbf{E}^{s}_{i})]_{i=0:t+1},[f_{k}(\mathbf{E}^{s}_{i})]_{i=0:t+1},[f_{v}(\mathbf{E}^{s}_{i})]_{i=0:t+1}),$ (25)
@@
$[\mathbf{u}^{r}_{i}]_{i=0:t}$ $=\text{Causal SelfAttn}(\mathbf{q},\mathbf{k},\mathbf{v})$
$=\text{Causal SelfAttn}([f_{q}(\mathbf{E}^{r}_{i})]_{i=0:t},[f_{k}(\mathbf{E}^{r}_{i})]_{i=0:t},[f_{v}(\mathbf{E}^{r}_{i})]_{i=0:t}),$ (27)
@@
$[h_{i}]_{i=1:t}$ $=\text{Causal CrossAttn}(\mathbf{q},\mathbf{k},\mathbf{v})$
$=\text{Causal CrossAttn}([f_{q}(\mathbf{u}^{s}_{i})]_{i=0:t+1},[f_{k}(\mathbf{u}^{s}_{i})]_{i=0:t+1},[f_{v}(\mathbf{u}^{r}_{i})]_{i=0:t}),$ (29)
@@
These representations $h_{t}$ are further compressed via a transformation parametrized by $\psi$ to a latent posterior distribution $\mathcal{N}(\mu_{\psi}(h_{t}),\sigma_{\psi}(h_{t}))$. All the previous stages describing the internal structure of the transformer encoder-decoder are visualized in Figure˜3(b).
@@ 3.3.3 Bayesian Inference and Evidence Lower-Bound
$\mathcal{L}_{\text{ELBO},t}=\mathbb{E}_{q_{\phi}(b_{t}|\tau_{0:t+1})}\bigl[$ $\mathcal{L}^{\text{S}}_{\text{Recon}}+\mathcal{L}^{\text{R}}_{\text{Recon}}+\mathcal{L}_{\text{Regularization}}\bigr]$ (37)
@@ 4.1 Context-Adaptive Meta-Reinforcement Learning
Context-adaptive meta-RL refers to a class of methods in which agents adapt to new tasks by conditioning the policy on an inferred latent representation from the context. Context is an umbrella term that might include past transitions or any task-related information available. Instead of relying solely on gradient-based parameter updates at test time, these methods learn a mapping from the raw information in trajectories to a latent task domain during training, which is then used for improved adaptation. The conditioned policy will ideally be enabled to make task-aware decisions. Prominent examples include PEARL [rakelly2019efficient], which infers a variational latent for conditioning an off-policy actor-critic agent.
Among the various context-adaptive strategies, state augmentation, technically concatenation, which embeds the learned representation directly into the observation space of the agent, stands out for its architectural simplicity and empirical robustness. As opposed to using hypernetworks [beck2023hypernetworks], which directly tune the parameters of policy networks according to the context representation, state augmentation allows the context-conditioned policy to retain the Markov property and scale more effectively with task complexity [beck2023survey]. State-augmented policies are particularly effective in continuous control environments and are compatible with both on-policy and off-policy training regimes. These methods strike a balance between inference flexibility and representational capacity, making them well-suited for meta-RL settings.
In addition to offering strong empirical results, context-adaptive approaches facilitate interpretability and modularity in policy design. By decoupling task inference from policy optimization, these methods enable more transparent analysis of task representations and provide a solution for incorporating structured priors under uncertainty [humplik2019taskinference]. As such, context-adaptive meta-RL forms a crucial foundation for more complex frameworks, including Bayes-adaptive meta-RL, which builds on the core idea of conditioning on task inference to achieve generalization.
@@ 4.2 Off-Policy Context-Adaptive Meta-RL
In this section, a few off-policy meta-RL frameworks will be discussed that use some version of context embedding in their models.
Rakelly et al. [rakelly2019efficient] proposed PEARL that uses a learned latent space $z$ to condition the meta-policy on each task. Therefore, the former policy $\pi_{\theta}(a\lvert s)$ will be redefined as $\pi_{\theta}(a\lvert s,z)$. Here, $z$ will be estimated by training a probabilistic model called inference network $q_{\phi}(z\lvert\mathbf{c})$, where $c$ is a context variable consisting of task-specific independently sampled transitions. Hence, for a task $\mathcal{T}_{\mathrm{i}}$, $\mathbf{c}$ is a set of transitions $\mathbf{c}^{\mathcal{T}_{\mathrm{i}}}_{1:N}$ where
@@
The main advantage of PEARL is its adaptability to off-policy training regimes with minimal storage overhead. In the presence of a new task $\mathcal{T}$ at inference time, the context $c^{\mathcal{T}}$ will be accumulated in the form of transitions during the collection of the new experiences and $q_{\phi}(z^{\mathcal{T}}\lvert\mathbf{c}^{\mathcal{T}})$ will estimate a latent vector corresponding to $\mathcal{T}$. $\pi_{\theta}(a\lvert s,z^{\mathcal{T}})$ will be conditioned to the acquired knowledge about the dynamics of $\mathcal{T}$ and make decisions accordingly. Although the introduction of $q_{\phi}(z\lvert\mathbf{c})$ will enable the automatic task encoding feature for this method, the downside of PEARL remains in disregarding the temporal structure of the context, which limits its application to hierarchical tasks and partially-observable MDPs.
Nam et al. [nam2022skill] present an extension to PEARL’s approach, SiMPL, that enables meta-learning on long-horizon, sparse-reward tasks using offline datasets. This approach involves extracting reusable skills and a skill prior from the offline data, meta-training a high-level policy to compose these skills into long-horizon behaviors, and rapidly adapting the meta-trained policy to solve unseen target tasks. Experimental results on continuous control tasks demonstrate a considerable advantage compared with PEARL [rakelly2019efficient].
Unlike PEARL, SiMPL uses offline transition datasets without rewards or task labels to extract short, reusable skills for solving unseen long-horizon tasks. A high-level policy, guided by a skill prior from offline data, explores the learned skill space, while a probabilistic encoder infers task information from context. The high-level policy selects latent skill vectors $z$, which condition the low-level policy to produce actions based on the current state and selected skill.
Another notable off-policy meta-RL method, TIGR [bing2023tigr], is a task-inference-based meta-RL method that adapts the Bayesian meta-RL formulation, regularly present in on-policy-based meta-RL algorithms, by introducing Gaussian variational autoencoders (VAE) and gated recurrent units (GRUs) to capture multi-modal task representations in its task inference model. While off-policy meta-RL methods, predominantly PEARL-based approaches, handle parametric adaptation quite well, they still struggle when faced with non-parametric variations in task distributions. These methods rely on structured priors and efficient latent space encodings but lack the flexibility to adapt to non-stationary and qualitatively distinct task variations.
@@ 4.3 Transformers in Meta-RL
Meta-learning approaches in Bayesian meta-RL can be categorized into recurrent models, gradient-based methods, and transformer-based architectures. Traditional methods such as RL2 use RNNs to encode past experiences into hidden states, allowing agents to infer task dynamics based on episodic memory. However, these approaches often suffer from information bottlenecks and fail to model long-range dependencies effectively. Note that this is usually the case for the BAMDP hyper-episodes of length $H^{+}$.
Gradient-based meta-learners, such as MAML, optimize model parameters to allow fast adaptation through a few gradient updates. While effective in certain settings, these methods are computationally expensive and struggle with hierarchical robotic tasks with complex temporal dependencies. Moreover, they typically rely on a fixed initialization, like MAML, which limits their ability to generalize effectively across highly diverse environments.
Recent advancements in transformer-based reinforcement learning have demonstrated how RL can be formulated as a sequence modeling problem, paving the way for architectures such as Decision Transformer (DT) [chen2021decision] and Online Decision Transformer (ODT) [zheng2022online]. Decision Transformer defines RL by casting it as a conditional sequence modeling task, leveraging autoregressive modeling to compensate for partial observability and output actions based on the history of interactions for desired future returns. This eliminates the need for explicit value functions or policy gradients, making RL more interpretable and stable. Expanding on this, ODT extends the capabilities of the Decision Transformer by adding online fine-tuning into the framework. These developments indicate how transformers might enhance memory-based meta-learning. Prompt-DT [xu2022prompting] evolves the original DT by adding a prompt token to label the task and providing an implicit MDP definition for few-shot generalization. The prompt used in Prompt-DT is a combination of several short demonstrations, each containing only a few timesteps. Although this design falls under the multi-task RL paradigm that takes advantage of external information provided for each task, it is a remarkable example of generalizable policy generation with transformers in RL.
Other transformer-based architectures, such as TrMRL [melo2022transformers] and HTrMRL [shala2024hierarchical], provide an alternative approach by leveraging self-attention mechanisms to process entire sequences in parallel. Unlike RNN-based methods, transformers do not require sequential processing, allowing for more efficient training. TrMRL introduces an episodic memory mechanism, refining task representations at multiple layers. The memory reinstatement mechanism in TrMRL reuses previously learned representations for faster adaptation in meta-learning tasks. The hierarchical structure of HTrMRL enables the agent to capture intra-episode and inter-episode relationships, facilitating mode detection and definition generalization across tasks.
Notably, a recent work, AMAGO [grigsby2023amago], is an off-policy transformer-based context-adaptive meta-RL algorithm. AMAGO uses long-sequence transformers, trained over full trajectories in parallel. They tried to address two major issues in training large transformer models, the memory bottlenecks and planning horizon limitation, validated in goal-conditioned tasks with sparse rewards. Like TrMRL and HTrMRL, they leverage the parallelism transformers for inter-episode, i.e., inter-task, generalization. While AMAGO combines sequence modeling and off-policy learning, its reliance on dense replay data and goal relabeling introduces unnecessary challenges in context-adaptive meta-RL. Moreover, AMAGO does not explicitly address task uncertainty with amortized inference. Compared to models that infer latent task distributions, AMAGO prioritizes direct goal adaptation with relabeling, offering a complementary parallel approach within transformer-based meta-RL.
Considering the mentioned body of work with transformers in meta-RL, transformers can improve upon previous architectures in Bayesian meta-reinforcement learning. Improvements may involve better uncertainty handling, capturing complex multi-scale dependencies, or enabling agents to generalize more efficiently. Compared to RNN-based models and gradient-based meta-learning techniques, transformer architectures such as TrMRL and HTrMRL hold promise for a more robust and scalable solution for temporally challenging and diverse robotic RL applications.

### v2

@@ Abstract
Reinforcement learning (RL) enables robots to operate in uncertain environments, but standard approaches often struggle with poor generalization to unseen tasks. Context-adaptive meta reinforcement learning addresses these limitations by conditioning on the task representation, yet they mostly rely on complete action information in the experience making task inference tightly coupled to a specific policy. This paper introduces Context Representation via Action-Free Transformer encoder–decoder (CRAFT), a belief model that infers task representations solely from sequences of states and rewards. By removing the dependence on actions, CRAFT decouples task inference from policy optimization, supports modular training, and leverages amortized variational inference for scalable belief updates. Built on a transformer encoder–decoder with rotary positional embeddings, the model captures long-range temporal dependencies and robustly encodes both parametric and non-parametric task variations. Experiments on the MetaWorld ML-10 robotic manipulation benchmark show that CRAFT achieves faster adaptation, improved generalization, and more effective exploration compared to context-adaptive meta-RL baselines. These findings highlight the potential of action-free inference as a foundation for scalable RL in robotic control. [^1] This work has been published in part in the doctoral dissertation of Amir M. Soufi Enayati [enayati2025practical].
@@ 1 Introduction
Although Bayes-optimal agents offer ideal exploration-exploitation trade-offs, they are intractable in complex domains. Posterior sampling and latent variable meta-learning attempt to approximate this behavior, encoding task uncertainty into structured representations to reduce reliance on extensive interactions. Yet, most meta-RL approaches struggle with significant task variability, often assuming smooth parametric changes between tasks, which limits their adaptability in real-world robotic scenarios involving non-parametric variations [zintgraf2019varibad]. A schematic of different approaches in context accumulation for exploration in meta-RL is depicted in Figure˜1.
Figure 1: Comparison of belief construction in meta-RL. Posterior sampling (left) infers a latent belief from trajectories to compress relevant information. Bayes-optimal agents (middle) update the full posterior over tasks using complete interaction histories, which is intractable in complex domains. Bayes-adaptive formulations (right) embed belief into the state with principled belief updates; the proposed method follows this formulation.
@@
To address current limitations, a framework is proposed in this paper for extracting long-term interepisodic features for Bayes-adaptive exploration. Additionally, CRAFT isolates the agent’s performance from task history by removing the condition of the belief model on the actions. This abstraction will allow learned behaviors to be reused and recombined in every stage of the training. Leveraging structured latent spaces and variational inference, the method facilitates strategic exploration and robust generalization, evaluated in a robotic simulation benchmark. This integration of probabilistic task embeddings and meta-learning techniques aims to ameliorate performance limitations and advance autonomy in RL-driven robot manipulation.
@@ 2.1 Meta-RL General Outline
MAML is a conventional baseline that many other implementations are based on. It is a gradient-based optimization problem to formulate proximity after a few adaptation steps. With the loss function for each task $\mathcal{T}_{\mathrm{i}}$ defined as
$\mathcal{L}_{\mathcal{T}_{\mathrm{i}}}(\pi_{\theta})=-\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta}}\left[\sum_{t=0}^{N_{\tau}}\gamma^{t}r_{t+1}\right],$ (1)
@@
Another challenge in meta-RL is the problem of generalization across tasks and domains. Meta-RL algorithms aim to learn meta-policies that can adapt to new tasks or environments with minimal additional training. However, achieving robust generalization remains a significant issue. Meta-learned policies often struggle to transfer knowledge to tasks or domains that differ significantly from the training distribution, resulting in poor performance and limited applicability. Yu et al. [yu2020meta] have shown that state-of-the-art meta-RL algorithms usually struggle in adaptation to an unseen robotic manipulation task after being exposed to ten or even 45 tasks in meta-training. This indicates that knowledge sharing among the training tasks is not easily generalizable to the test domain.
@@ 2.2 Bayesian Approach To Context-Adaptive Meta-RL
$=\mathbb{E}_{b_{t}}\left[\mathcal{P}_{S}(s_{t+1}|s_{t},a_{t+1})\right]~\delta\Bigl(b_{t+1}=p(\mathcal{P}_{R},\mathcal{P}_{S}|\tau_{0:t+1})\Bigl).$ (6)
@@ 3.2 The Action-Free Belief Model Design
$\mathcal{P}_{R}(r|s,a,s^{\prime})\approx\tilde{\mathcal{P}}_{R}(r|s_{0},\dots,s_{-1},s,s^{\prime}),$
and,
$\mathcal{P}_{S}(s^{\prime}|s,a)\approx\tilde{\mathcal{P}}_{S}(s^{\prime}|s_{0},\dots,s_{-1},s).$
@@ 3.3.1 Rotary Positional Embedding
where $i$ is the token position in the context and $t\in{0,\dots,d/2-1}$ with $d$ representing the length of the token.
@@ 3.3.2 Implementation in CRAFT
The architecture shown in Figure˜2(a) uses a transformer encoder-decoder to extract compact and informative representations from action-free trajectories, aligning with the structure of belief updates in Bayes-Adaptive Markov Decision Processes (BAMDP). Borrowing from Equation˜12, the posterior belief over task identity is defined as
@@
[…]
=Causal SelfAttn($\text${Causal SelfAttn}( $[f_{q}(\mathbf{E}^{s}_{i})]_{i=0:t+1},$
$[f_{k}(\mathbf{E}^{s}_{i})]_{i=0:t+1},$
$[f_{v}(\mathbf{E}^{s}_{i})]_{i=0:t+1}),$ (25)
@@
[…]
=Causal SelfAttn($\text${Causal SelfAttn}( $[f_{q}(\mathbf{E}^{r}_{i})]_{i=0:t},$
$[f_{k}(\mathbf{E}^{r}_{i})]_{i=0:t},$
$[f_{v}(\mathbf{E}^{r}_{i})]_{i=0:t}),$ (27)
@@
[…]
=Causal CrossAttn($\text${Causal CrossAttn}( $[f_{q}(\mathbf{u}^{s}_{i})]_{i=0:t+1},$
$[f_{k}(\mathbf{u}^{s}_{i})]_{i=0:t+1},$
$[f_{v}(\mathbf{u}^{r}_{i})]_{i=0:t}),$ (29)
@@
These representations $h_{t}$ are further compressed via a transformation parametrized by $\psi$ to a latent posterior distribution $\mathcal{N}(\mu_{\psi}(h_{t}),\sigma_{\psi}(h_{t}))$. All the previous stages describing the internal structure of the transformer encoder-decoder are visualized in Figure˜2(b).
@@ 3.3.3 Bayesian Inference and Evidence Lower-Bound
$\mathcal{L}_{\text{ELBO},t}=\mathbb{E}_{q_{\phi}(b_{t}|\tau_{0:t+1})}\bigl[$ $\mathcal{L}^{\text{S}}_{\text{Recon}}+\mathcal{L}^{\text{R}}_{\text{Recon}}+\mathcal{L}_{\text{Regularization}}\bigr]$
@@ 4.1 Context-Adaptive Meta-Reinforcement Learning
Context-adaptive meta-RL enables agents to adapt to new tasks by conditioning policies on an inferred latent representation of the context, which may include transitions or other task-related information. Rather than performing gradient-based updates at test time, these methods learn a mapping from trajectory data to a latent task embedding used for fast adaptation. PEARL [rakelly2019efficient] is a representative framework that infers a variational latent for conditioning an off-policy actor–critic agent.
Among existing strategies, state augmentation, i.e., concatenating the learned representation with the observation, offers simplicity and robustness. Unlike hypernetwork-based approaches [beck2023hypernetworks], which tune policy parameters directly from the context, state augmentation preserves the Markov property and scales effectively with task complexity [beck2023survey]. These context-conditioned policies perform well in continuous control and integrate smoothly with both on- and off-policy methods. Moreover, separating inference from policy optimization improves interpretability and modularity, allowing structured priors to be incorporated under uncertainty [humplik2019taskinference]. Such context-adaptive designs form the foundation for more advanced Bayes-adaptive meta-RL frameworks that build on explicit task inference for generalization.
@@ 4.2 Off-Policy Context-Adaptive Meta-RL
Several off-policy meta-RL algorithms incorporate context embeddings. PEARL [rakelly2019efficient] learns a latent variable $z$ to condition the policy $\pi_{\theta}(a|s,z)$ using an inference network $q_{\phi}(z|\mathbf{c})$, where $\mathbf{c}$ is a set of task-specific transitions,
@@
At inference, $q_{\phi}(z^{\mathcal{T}}|\mathbf{c}^{\mathcal{T}})$ updates as new transitions are observed, allowing $\pi_{\theta}$ to adapt rapidly. Its main drawback lies in neglecting temporal structure, limiting application to hierarchical or partially observable domains.
SiMPL [nam2022skill] extends PEARL for long-horizon, sparse-reward tasks using offline data. It extracts reusable skills and a skill prior, then meta-trains a high-level policy to compose these skills into longer behaviors. The high-level policy samples latent skill vectors $z$ conditioning a low-level controller, enabling adaptation without reward or task labels. Experiments on continuous control tasks show significant gains over PEARL.
TIGR [bing2023tigr] further generalizes Bayesian meta-RL to the off-policy setting using Gaussian variational autoencoders and gated recurrent units to capture multi-modal task representations. Although such PEARL-based methods handle parametric variations effectively, they struggle with non-parametric or non-stationary task distributions, relying on structured priors and fixed latent spaces that limit flexibility.
@@ 4.3 Transformers in Meta-RL
Meta-learning algorithms for Bayesian meta-RL include recurrent, gradient-based, and transformer-based models. Recurrent methods like RL2 encode experience through hidden states but suffer from information bottlenecks over long meta-horizons $H^{+}$. Gradient-based learners such as MAML enable fast adaptation via a few gradient steps but are computationally heavy and often fail to generalize across diverse or hierarchical tasks.
Transformers reframe RL as sequence modeling. Decision Transformer (DT) [chen2021decision] models trajectories autoregressively to output actions conditioned on desired returns, removing the need for explicit value functions. Online Decision Transformer (ODT) [zheng2022online] adds online fine-tuning, and Prompt-DT [xu2022prompting] incorporates short demonstration prompts as task labels for few-shot generalization. Though positioned within multi-task RL, Prompt-DT exemplifies transformer-driven generalizable policy generation.
Architectures such as TrMRL [melo2022transformers] and HTrMRL [shala2024hierarchical] apply self-attention for parallel sequence processing, improving efficiency and long-term memory. TrMRL introduces episodic memory reinstatement for faster adaptation, while HTrMRL captures hierarchical dependencies across and within episodes. AMAGO [grigsby2023amago] combines off-policy learning with long-sequence transformers to mitigate memory and horizon limitations in sparse-reward, goal-conditioned tasks. Although effective for large-scale parallel training, AMAGO does not explicitly address task uncertainty through amortized inference, focusing instead on goal relabeling.
Overall, transformer-based meta-RL improves scalability, long-range credit assignment, and uncertainty modeling compared to recurrent and gradient-based methods. Architectures such as TrMRL, HTrMRL, and AMAGO demonstrate the growing potential of sequence models for efficient adaptation and generalization in robotic RL.